# DeformTime: Capturing Variable Dependencies with Deformable Attention for Time Series Forecasting

### v2

@@ 3.4 Temporal deformable attention block (T-DAB)
The factor $r$ (time steps) can be set to different values across encoder layers to capture dependencies at various temporal granularities (see Appendix C). If we set $r=1$, $\mathbf{Z}_{\texttt{e}}$ remains unchanged. Similarly to V-DAB, we obtain the query embedding $\mathbf{Q}_{\texttt{r}}\!\in\!\mathbb{R}^{\kappa\times d}$ over the transformed input using $\mathbf{Q}_{\texttt{r}}=\mathbf{Z}_{\texttt{r}}\mathbf{U}_{Q}$, where $\mathbf{U}_{Q}\!\in\!\mathbb{R}^{d\times d}$ is a weight matrix. We then implement an augmented Transformer attention that contains deformed information across the temporal dimension. We work our way column-wise, and for every column $\mathbf{z}_{\texttt{r}}\!\in\!\mathbf{Z}_{\texttt{r}}$, we sample (linearly as opposed to the bilinear approach in V-DAB) from an index position $p+\Delta p\!\in\!\mathbb{R}$.
@@
where $\phi(\cdot)$ denotes a linear interpolation function over $2$ adjacent points of $\mathbf{z}_{\texttt{r}}\!\in\!\mathbf{Z}_{\texttt{r}}^{(g)}$ determined by the deformed index $p+\Delta p^{(g)}$.
Similarly to V-DAB, $\mathbf{K}_{\texttt{s}},\mathbf{V}_{\texttt{s}}\!\in\!\mathbb{R}^{\kappa\times d}$ are obtained from $\mathbf{Z}_{\texttt{s}}$, using $\mathbf{K}_{\texttt{s}}=\mathbf{Z}_{\texttt{s}}\mathbf{U}_{K}$, and $\mathbf{V}_{\texttt{s}}=\mathbf{Z}_{\texttt{s}}\mathbf{U}_{V}+\mathbf{P}_{\texttt{t}}$, where $\mathbf{P}_{\texttt{t}}\!\in\!\mathbb{R}^{\kappa\times d}$ is a relative positional bias. We then apply multi-head attention, by first splitting $\mathbf{K}_{\texttt{s}}$ and $\mathbf{V}_{\texttt{s}}$ into $G$ groups (submatrices) denoted by $\mathbf{K}_{g}$ and $\mathbf{V}_{g}\!\in\!\mathbb{R}^{\kappa\times(d/G)}$, and then by using $\mathbf{A}_{g}=\texttt{softmax}\!\left(\mathbf{Q}_{g}\mathbf{K}_{g}^{\top}/\sqrt{d/G}\right)\mathbf{V}_{g}$. The outputs from different heads are concatenated column-wise and linearly projected with a weight matrix $\mathbf{W}_{i}\!\in\!\mathbb{R}^{d\times d}$ to obtain the attention embedding of the $i$-th patch $\mathbf{A}_{i}\!\in\!\mathbb{R}^{\kappa\times d}$. To form the output of T-DAB, denoted by $\mathbf{Z}_{\texttt{t}}\!\in\!\mathbb{R}^{L\times d}$, the outputs of all patches are concatenated along the first dimension and re-arranged back to the original temporal structure (matching the input $\mathbf{Z}_{\texttt{e}}$ as well as V-DAB’s output $\mathbf{Z}_{\texttt{v}}$).
@@ 3.5 Encoder
The encoder of DeformTime may consist of more than one layer. In our experiments, we deploy $2$ encoder layers, similarly to related work [28, 46]. Each encoder layer contains two DAB branches (on the left and right side, respectively – see also Figure 1), comprising stacked transformer blocks [8] with layer normalisation placed within the residual connection [49, 48]. Instead of using vanilla attention, we are introducing a V-DAB (left side) and a T-DAB (right side) module to capture inter- and intra-variate dependencies, respectively. The operations of a branch can be summarised by
$\mathbf{Z}_{\texttt{i}}=\texttt{Drop}\!\left(\texttt{DAB}\left(\mathbf{Z}_{\texttt{e}}\right)\right)+\mathbf{Z}_{\texttt{e}}\,\,\text{followed by}\,\,\mathbf{Z}_{\texttt{c}}=\texttt{Drop}\!\left(\texttt{MLP}\!\left(\texttt{LN}\!\left(\mathbf{Z}_{\texttt{i}}\right)\right)\right)+\mathbf{Z}_{\texttt{e}}\,,$ (10)
where $\texttt{DAB}(\cdot)$ denotes the V-DAB or T-DAB operation, $\mathbf{Z}_{\texttt{i}},\mathbf{Z}_{\texttt{c}}\!\in\!\mathbb{R}^{L\times d}$ are an intermediate output and the output of the DAB encoder layer, respectively, $\texttt{Drop}(\cdot)$ denotes stochastic depth [12] where layers within the network are randomly dropped during training with a learnable probability, and $\texttt{MLP}(\cdot)$ is a $2$-layer perceptron with ReLU activation and a hidden size of $d$. The outputs from the two encoder branches are concatenated over the second dimension. We feed the concatenated output into a fully connected layer to form the output $\mathbf{Z}_{j}\!\in\!\mathbb{R}^{L\times d}$ of the $j$-th encoder layer.
@@ 3.6 Encoder-decoder structure
In T-DAB (section 3.4), we set the time window $r$ to different values across encoder layers, enabling the model to attend to information at multiple temporal granularities. To effectively learn from this encoding, we construct a hierarchical encoder structure motivated by prior related work [57, 26, 55]. Specifically, the output from encoder layer $j\!-\!1$, $\mathbf{Z}_{j-1}\!\in\!\mathbb{R}^{L\times d}$, becomes the input to encoder layer $j$. This is formulated by
@@ 4 Results
We present forecasting accuracy results across $3$ benchmark and $3$ disease rate prediction tasks, comparing DeformTime to other competitive ‘baseline’ models. In addition, we provide an ablation study and a computational complexity / efficiency analysis.
4.1 Forecasting tasks, baseline methods, and experiment settings
@@
Experiments are conducted on $6$ real-world data sets. These include $3$ established benchmarks from previously published papers [45, 53, 52, 46, 32, 29], namely a weather and $2$ electricity transformer temperature (ETTh1 and ETTh2) tasks with respective temporal resolutions of $10$ minutes and $1$ hour. In addition, we have formed $3$ disease rate modelling tasks, focusing on the prediction of influenza-like illness (ILI) rates in England (ILI-ENG) and US Health & Human Services (HHS) Regions 2 and 9 (ILI-US2 and ILI-US9). For the ILI rate forecasting tasks, we also introduce the frequency time series of web searches as exogenous predictors. In the ETT and weather tasks, we consider oil and wet-bulb temperature, respectively, as the target variables; we refer to the rest indicators as exogenous variables. We note that the ILI forecasting tasks have considerably more exogenous variables compared to the ETTh1, ETTh2, and weather tasks. Based on the evaluation settings in prior work [32, 28], for the ETT and weather tasks, we set the forecasting horizon $H$ to $96$, $192$, $336$, and $720$ time steps, and use a single test fold of consecutive unseen instances. For the ILI forecasting task, we conduct experiments on $4$ consecutive influenza seasons (2015/16 to 2018/19) as separate test sets ($4$ test folds); the forecasting horizons we consider are $H\!=\!\{7,14,21,28\}$ days ahead. Further details, including the evaluation setup are available in Appendix A.
@@
We compare DeformTime to 6 competitive forecasting models that were initially designed for MTS tasks and, to the best of our knowledge, form the current SOTA methods. These are LightTS [54], DLinear [53], Crossformer [55], PatchTST [32], iTransformer [28], and TimeMixer [44]. We also include a naïve persistence model baseline (Appendix B). For the ILI forecasting tasks, we conduct hyperparameter tuning for all models. In the benchmark tasks, with the exception of Crossformer and LightTS, we adopt the settings from their official repositories to reproduce results. Further details about the experiment settings can be found in Appendix C.
@@
Table 1: Forecasting accuracy results across all tasks and methods. $H$ denotes the forecasting horizon time steps. For the ILI forecasting tasks, the table enumerates the average error across the $4$ test seasons. Complete results per season are shown in Appendix D. $\epsilon\leavevmode\nobreak\ \%$ denotes sMAPE (to reduce column width). The best results are in bold font and the second best are underlined.
@@
Models DeformTime PatchTST iTransformer TimeMixer Crossformer LightTS DLinear Persistence
$H$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$
ETTh1 96 0.1941 14.96 0.2017 15.41 0.2052 15.46 0.2112 16.32 0.2126 16.52 0.2215 17.24 0.2599 20.82 0.2371 18.47
192 0.2116 16.08 0.2409 18.29 0.2429 18.13 0.2382 17.91 0.2820 21.63 0.2636 20.55 0.3798 31.78 0.2803 21.46
336 0.2158 16.27 0.2559 19.29 0.2593 19.11 0.2625 19.72 0.2947 22.65 0.2807 22.15 0.6328 58.34 0.3028 22.90
720 0.2862 21.81 0.3087 23.89 0.2886 22.05 0.3055 23.25 0.3350 24.84 0.5334 44.57 0.7563 69.52 0.3222 25.29
ETTh2 96 0.3121 40.07 0.3145 39.25 0.3420 42.41 0.3454 41.27 0.3486 40.71 0.3507 41.80 0.3349 41.68 0.3522 43.85
192 0.3281 37.90 0.3839 45.45 0.4233 47.44 0.4183 47.49 0.4035 43.16 0.4022 48.01 0.4084 50.67 0.4416 50.24
336 0.3450 37.00 0.4018 46.77 0.4332 45.95 0.4380 46.79 0.4487 49.44 0.4425 51.35 0.4710 55.53 0.4836 53.70
720 0.3640 34.99 0.4960 55.27 0.4565 45.40 0.4729 46.37 0.5832 61.45 0.6252 70.50 0.7981 94.67 0.5199 58.75
Weather 96 0.0244 37.89 0.0258 39.37 0.0277 42.39 0.0322 45.90 0.0271 44.92 0.0293 48.48 0.0251 39.03 0.0329 51.83
192 0.0260 39.33 0.0279 42.02 0.0277 42.77 0.0347 48.62 0.0308 54.14 0.0319 51.45 0.0270 42.68 0.0361 54.92
336 0.0291 44.26 0.0303 45.31 0.0308 46.01 0.0359 49.75 0.0345 62.53 0.0317 50.83 0.0305 47.68 0.0361 55.14
720 0.0363 53.72 0.0389 56.04 0.0395 57.01 0.0457 59.82 0.0395 65.47 0.0386 62.96 0.0352 54.54 0.0394 56.04
ILI-ENG 7 1.6417 28.61 2.3115 27.61 2.3084 26.38 2.1748 25.68 1.8698 25.71 2.2397 52.25 2.8214 43.02 2.1710 24.96
14 2.2308 33.98 3.2547 37.76 3.2301 36.67 3.0209 35.39 2.6543 30.97 2.6879 38.29 3.7922 55.28 3.0625 33.77
21 2.6500 32.70 4.3192 51.11 4.2347 48.93 3.5501 49.36 3.0014 40.57 3.3616 51.78 4.4739 61.25 3.8617 42.03
28 2.7228 40.44 4.9964 59.60 4.8125 55.35 4.1188 54.60 3.1983 46.14 3.4132 55.59 5.0347 67.75 4.5857 49.49
ILI-US2 7 0.4122 16.01 0.7097 24.52 0.6507 23.24 0.5284 20.07 0.4400 16.46 0.4632 16.74 0.7355 27.94 0.6474 22.48
14 0.4752 17.73 0.8635 30.11 0.7896 28.17 0.6556 24.61 0.5852 20.98 0.5827 23.11 0.8435 32.22 0.8135 28.24
21 0.5425 22.13 1.0286 36.70 0.8042 30.03 0.6794 27.68 0.6245 22.29 0.6683 29.27 0.9124 34.93 0.9635 33.51
28 0.5538 22.25 1.1525 42.61 0.9619 36.75 0.8853 36.53 0.6512 23.91 0.7175 27.73 0.9805 37.62 1.1007 38.54
ILI-US9 7 0.2622 12.26 0.4116 19.34 0.4057 18.57 0.3239 15.21 0.3149 14.44 0.3185 15.65 0.4675 23.47 0.4057 18.49
14 0.3084 13.80 0.5020 24.09 0.4702 22.44 0.4060 19.08 0.3571 17.23 0.3791 19.04 0.5467 27.35 0.5008 23.07
21 0.3179 14.24 0.5935 29.40 0.5106 24.11 0.4576 21.40 0.3418 15.90 0.4754 23.74 0.6001 29.66 0.5906 27.41
28 0.3532 15.74 0.6665 33.35 0.6498 31.04 0.5124 24.11 0.3747 16.44 0.4769 23.22 0.6564 32.16 0.6799 31.67

### v3

@@ 3.4 Temporal deformable attention block (T-DAB)
The factor $r$ (time steps) can be set to different values across encoder layers to capture dependencies at various temporal granularities (see Appendix D). If we set $r=1$, $\mathbf{Z}_{\texttt{e}}$ remains unchanged. Similarly to V-DAB, we obtain the query embedding $\mathbf{Q}_{\texttt{r}}\!\in\!\mathbb{R}^{\kappa\times d}$ over the transformed input using $\mathbf{Q}_{\texttt{r}}=\mathbf{Z}_{\texttt{r}}\mathbf{U}_{Q}$, where $\mathbf{U}_{Q}\!\in\!\mathbb{R}^{d\times d}$ is a weight matrix. We then implement an augmented Transformer attention that contains deformed information across the temporal dimension. We work our way column-wise, and for every column $\mathbf{z}_{\texttt{r}}\!\in\!\mathbf{Z}_{\texttt{r}}$, we sample (linearly as opposed to the bilinear approach in V-DAB) from an index position $p+\Delta p\!\in\!\mathbb{R}$.
@@
where $\phi(\cdot)$ denotes a linear interpolation function over $2$ adjacent points of $\mathbf{z}_{\texttt{r}}\!\in\!\mathbf{Z}_{\texttt{r}}^{(g)}$ determined by the deformed index $p+\Delta p^{(g)}$. If a sampling position is outside of the matrix boundaries, we interpolate with 00.
Similarly to V-DAB, $\mathbf{K}_{\texttt{s}},\mathbf{V}_{\texttt{s}}\!\in\!\mathbb{R}^{\kappa\times d}$ are obtained from $\mathbf{Z}_{\texttt{s}}$, using $\mathbf{K}_{\texttt{s}}=\mathbf{Z}_{\texttt{s}}\mathbf{U}_{K}$, and $\mathbf{V}_{\texttt{s}}=\mathbf{Z}_{\texttt{s}}\mathbf{U}_{V}+\mathbf{P}_{\texttt{t}}$, where $\mathbf{P}_{\texttt{t}}\!\in\!\mathbb{R}^{\kappa\times d}$ is a relative positional bias. $\mathbf{K}_{\texttt{s}}$ and $\mathbf{V}_{\texttt{s}}$ thus contain elements sampled from the neighbouring time steps $(z_{i,j-\alpha},z_{i,j+\alpha})$. We then apply multi-head attention to have elements at the position $p$ attend to elements at the deformed position $p+\Delta p^{(g)}$. This is achieved by first splitting $\mathbf{K}_{\texttt{s}}$ and $\mathbf{V}_{\texttt{s}}$ into $G$ groups (submatrices) denoted by $\mathbf{K}_{g}$ and $\mathbf{V}_{g}\!\in\!\mathbb{R}^{\kappa\times(d/G)}$, and then by using $\mathbf{A}_{g}=\texttt{softmax}\!\left(\mathbf{Q}_{g}\mathbf{K}_{g}^{\top}/\sqrt{d/G}\right)\mathbf{V}_{g}$. The outputs from different heads are concatenated column-wise and linearly projected with a weight matrix $\mathbf{W}_{i}\!\in\!\mathbb{R}^{d\times d}$ to obtain the attention embedding of the $i$-th patch $\mathbf{A}_{i}\!\in\!\mathbb{R}^{\kappa\times d}$. To form the output of T-DAB, denoted by $\mathbf{Z}_{\texttt{t}}\!\in\!\mathbb{R}^{L\times d}$, the outputs of all patches are concatenated along the first dimension and re-arranged back to the original temporal structure (matching the input $\mathbf{Z}_{\texttt{e}}$ as well as V-DAB’s output $\mathbf{Z}_{\texttt{v}}$).
@@ 3.5 Encoder
The encoder of DeformTime may consist of more than one layer. In our experiments, we deploy $2$ encoder layers, similarly to related work (Liu et al., 2024; Wu et al., 2023). Each encoder layer contains two DAB branches (on the left and right side, respectively – see also Figure 1), comprising stacked transformer blocks (Dosovitskiy et al., 2021) with layer normalisation placed within the residual connection (Xiong et al., 2020; Xia et al., 2022). Instead of using vanilla attention, we are introducing a V-DAB (left side) and a T-DAB (right side) module to capture inter- and intra-variable dependencies, respectively. The operations of a branch can be summarised by
$\mathbf{Z}_{\texttt{i}}$ $=\texttt{Drop}\!\left(\texttt{DAB}\left(\mathbf{Z}_{\texttt{e}}\right)\right)+\mathbf{Z}_{\texttt{e}}\,\,\,\,\text{and}$ (10)
$\mathbf{Z}_{\texttt{c}}$ $=\texttt{Drop}\!\left(\texttt{MLP}\!\left(\texttt{LN}\!\left(\mathbf{Z}_{\texttt{i}}\right)\right)\right)+\mathbf{Z}_{\texttt{e}}\,,$ (11)
where $\texttt{DAB}(\cdot)$ denotes the V-DAB or T-DAB operation, $\mathbf{Z}_{\texttt{i}},\mathbf{Z}_{\texttt{c}}\!\in\!\mathbb{R}^{L\times d}$ are an intermediate output and the output of the DAB encoder layer, respectively, $\texttt{Drop}(\cdot)$ denotes stochastic depth (Huang et al., 2016) where layers within the network are randomly dropped during training with a learnable probability, and $\texttt{MLP}(\cdot)$ is a $2$-layer perceptron with ReLU activation and a hidden size of $d$. The outputs from the two encoder branches are concatenated over the second dimension. We feed the concatenated output into a fully connected layer to form the output $\mathbf{Z}_{j}\!\in\!\mathbb{R}^{L\times d}$ of the $j$-th encoder layer.
@@ 3.6 Encoder-decoder structure
In T-DAB (section 3.4), we set the time window $r$ to different values across encoder layers, enabling the model to attend to information at multiple temporal granularities. To effectively learn from this encoding, we construct a hierarchical encoder structure motivated by prior related work (Zhou et al., 2021; Liu et al., 2022a; Zhang & Yan, 2023). Specifically, the output from encoder layer $j\!-\!1$, $\mathbf{Z}_{j-1}\!\in\!\mathbb{R}^{L\times d}$, becomes the input to encoder layer $j$. This is formulated by
@@ 4 Results
We present forecasting accuracy results across $3$ benchmark and $3$ disease rate prediction tasks, comparing DeformTime to $9$ competitive baseline models. In addition, we provide an ablation study and a computational complexity / efficiency analysis.
4.1 Experimental settings
Forecasting tasks and baseline methods.
@@ 4.1 Experimental settings
Experiments are conducted on $6$ real-world data sets. These include $3$ established benchmarks from previously published papers (Wu et al., 2021; Zeng et al., 2023; Yi et al., 2023b; Wu et al., 2023; Nie et al., 2023; Luo & Wang, 2024b), and specifically a weather-related and $2$ electricity transformer temperature (ETTh1 and ETTh2) forecasting tasks with respective temporal resolutions of $10$ minutes and $1$ hour. In addition, we have formed $3$ disease prevalence modelling tasks, focusing on the prediction of influenza-like illness (ILI) rates in England (ILI-ENG) and US Health & Human Services (HHS) Regions 2 and 9 (ILI-US2 and ILI-US9). For the ILI rate forecasting tasks, we also introduce the frequency time series of web searches as exogenous predictors. In the ETT and weather tasks, we consider oil temperature and carbon dioxide level, respectively, as the target variables; we refer to the rest indicators as exogenous variables. We note that the ILI forecasting tasks have considerably more exogenous variables compared to the ETTh1, ETTh2, and weather tasks. Based on the evaluation settings in prior work (Nie et al., 2023; Liu et al., 2024), for the ETT and weather tasks, we set the forecasting horizon $H$ to $96$, $192$, $336$, and $720$ time steps ahead, and use a single test fold of consecutive unseen instances. For the ILI forecasting task, we conduct experiments on $4$ consecutive influenza seasons (2015/16 to 2018/19) as separate test sets ($4$ test folds); the forecasting horizons we consider are $H\!=\!\{7,14,21,28\}$ days ahead. Further details, including the evaluation setup are available in Appendix A. While we are aware of other datasets used in MTS forecasting, not all of them are suitable because they may not have an explicitly defined target variable or pursue an ill-defined prediction task (see Appendix A.4 for further justification).
@@
We compare DeformTime to $9$ competitive time series forecasting models that to the best of our knowledge, form the current SOTA methods. These are LightTS (Zhang et al., 2022), DLinear (Zeng et al., 2023), Crossformer (Zhang & Yan, 2023), PatchTST (Nie et al., 2023), iTransformer (Liu et al., 2024), TimeMixer (Wang et al., 2024a), ModernTCN (Luo & Wang, 2024b), CycleNet (Lin et al., 2024), and TimeXer (Wang et al., 2024b). We also include a naïve persistence model baseline. Further details and justification for the selection of baselines can be found in Appendix C. For the ILI forecasting tasks, we conduct hyperparameter tuning for all models. In the other benchmark tasks, with the exception of Crossformer and LightTS, we adopt the settings from their official repositories to reproduce results.
Note that due to different task formulations, DeformTime and TimeXer focus on forecasting the endogenous variables only, whereas the other baseline models also predict future values of the variables that we consider as exogenous. Notably, we only use the last value of the predicted target variable $y_{t+H}$ (see section 2) for model evaluation, regardless of the number of outputs each model makes (see also Appendix D).
@@
Table 1: Forecasting accuracy results across all tasks and methods. $H$ denotes the forecasting horizon time steps. For the ILI forecasting tasks, the table enumerates the average error across the $4$ test seasons. Complete results per season are shown in Appendix E. $\epsilon\leavevmode\nobreak\ \%$ denotes sMAPE. The best results are in bold font and the second best are underlined.
@@
Models DeformTime ModernTCN CycleNet TimeXer PatchTST iTransformer TimeMixer Crossformer LightTS DLinear Persistence
$H$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$ MAE $\epsilon\leavevmode\nobreak\ \%$
ETTh1 96 0.1941 14.96 0.2047 15.66 0.1976 15.05 0.2135 16.03 0.2017 15.41 0.2052 15.46 0.2112 16.32 0.2126 16.52 0.2215 17.24 0.2599 20.82 0.2371 18.47
192 0.2116 16.08 0.2417 18.32 0.2411 18.25 0.2322 16.96 0.2409 18.29 0.2429 18.13 0.2382 17.91 0.2820 21.63 0.2636 20.55 0.3798 31.78 0.2803 21.46
336 0.2158 16.27 0.2415 18.52 0.2425 18.35 0.2414 17.34 0.2559 19.29 0.2593 19.11 0.2625 19.72 0.2947 22.65 0.2807 22.15 0.6328 58.34 0.3028 22.90
720 0.2862 21.81 0.2785 20.44 0.3018 22.98 0.2617 18.58 0.3087 23.89 0.2886 22.05 0.3055 23.25 0.3350 24.84 0.5334 44.57 0.7563 69.52 0.3222 25.29
ETTh2 96 0.3121 40.07 0.3199 40.68 0.3282 40.74 0.3346 41.04 0.3145 39.25 0.3420 42.41 0.3454 41.27 0.3486 40.71 0.3507 41.80 0.3349 41.68 0.3522 43.85
192 0.3281 37.90 0.3887 47.08 0.3687 41.78 0.4154 47.07 0.3839 45.45 0.4233 47.44 0.4183 47.49 0.4035 43.16 0.4022 48.01 0.4084 50.67 0.4416 50.24
336 0.3450 37.00 0.3904 50.54 0.3815 42.57 0.4041 42.26 0.4018 46.77 0.4332 45.95 0.4380 46.79 0.4487 49.44 0.4425 51.35 0.4710 55.53 0.4836 53.70
720 0.3640 34.99 0.5728 63.04 0.4827 49.93 0.5135 56.17 0.4960 55.27 0.4565 45.40 0.4729 46.37 0.5832 61.45 0.6252 70.50 0.7981 94.67 0.5199 58.75
Weather 96 0.0244 37.89 0.0279 42.49 0.0253 38.87 0.0301 44.04 0.0258 39.37 0.0277 42.39 0.0322 45.90 0.0271 44.92 0.0293 48.48 0.0251 39.03 0.0329 51.83
192 0.0260 39.33 0.0314 46.05 0.0270 41.72 0.0326 46.85 0.0279 42.02 0.0277 42.77 0.0347 48.62 0.0308 54.14 0.0319 51.45 0.0270 42.68 0.0361 54.92
336 0.0291 44.26 0.0351 50.12 0.0301 45.31 0.0348 49.12 0.0303 45.31 0.0308 46.01 0.0359 49.75 0.0345 62.53 0.0317 50.83 0.0305 47.68 0.0361 55.14
720 0.0363 53.72 0.0414 57.66 0.0383 56.47 0.0437 58.70 0.0389 56.04 0.0395 57.01 0.0457 59.82 0.0395 65.47 0.0386 62.96 0.0352 54.54 0.0394 56.04
ILI-ENG 7 1.6417 28.61 1.9489 28.28 2.5554 29.59 2.8084 33.66 2.3115 27.61 2.3084 26.38 2.1748 25.68 1.8698 25.71 2.2397 52.25 2.8214 43.02 2.1710 24.96
14 2.2308 33.98 2.7050 36.01 3.3911 39.42 3.4937 41.88 3.2547 37.76 3.2301 36.67 3.0209 35.39 2.6543 30.97 2.6879 38.29 3.7922 55.28 3.0625 33.77
21 2.6500 32.70 3.0400 40.02 4.4519 52.83 4.3337 51.56 4.3192 51.11 4.2347 48.93 3.5501 49.36 3.0014 40.57 3.3616 51.78 4.4739 61.25 3.8617 42.03
28 2.7228 40.44 3.3611 47.87 5.0259 59.93 4.9013 61.60 4.9964 59.60 4.8125 55.35 4.0000 54.27 3.1983 46.14 3.4132 55.59 5.0347 67.75 4.5857 49.49
ILI-US2 7 0.4122 16.01 0.4398 16.55 0.6951 24.67 0.6083 23.38 0.7097 24.52 0.6507 23.24 0.5284 20.07 0.4400 16.46 0.4632 16.74 0.7355 27.94 0.6474 22.48
14 0.4752 17.73 0.5279 20.22 0.8219 29.15 0.7725 29.07 0.8635 30.11 0.7896 28.17 0.6556 24.61 0.5852 20.98 0.5827 23.11 0.8435 32.22 0.8135 28.24
21 0.5425 22.13 0.5781 23.85 1.0469 37.98 0.8243 31.46 1.0286 36.70 0.8042 30.03 0.6794 27.68 0.6245 22.29 0.6683 29.27 0.9124 34.93 0.9635 33.51
28 0.5538 22.25 0.5710 23.66 1.1388 42.31 0.9074 34.72 1.1525 42.61 0.9619 36.75 0.8853 36.53 0.6512 23.91 0.7175 27.73 0.9805 37.62 1.1007 38.54
ILI-US9 7 0.2622 12.26 0.2899 14.17 0.4480 21.05 0.3813 18.20 0.4116 19.34 0.4057 18.57 0.3239 15.21 0.3149 14.44 0.3185 15.65 0.4675 23.47 0.4057 18.49
14 0.3084 13.80 0.3417 15.29 0.5072 24.02 0.4665 22.14 0.5020 24.09 0.4702 22.44 0.4060 19.08 0.3571 17.23 0.3791 19.04 0.5467 27.35 0.5008 23.07
21 0.3179 14.24 0.3710 15.43 0.5926 28.47 0.5715 27.42 0.5935 29.40 0.5106 24.11 0.4576 21.40 0.3418 15.90 0.4754 23.74 0.6001 29.66 0.5906 27.41
28 0.3532 15.74 0.3940 17.19 0.7031 34.50 0.6555 31.32 0.6665 33.35 0.6498 31.04 0.5124 24.11 0.3747 16.44 0.4769 23.22 0.6564 32.16 0.6799 31.67
DeformTime’s setup.
We use $G\!=\!4$ variable neighbourhoods in all our experiments (see section 3.2). This setting ($G\!=\!4$) also determines the number of group partitions for offset generation and multi-head attention in T-DAB (section 3.4). We provide a more detailed analysis of the effect of this specific parameter in Appendix E.4. The learnable amplitude $\alpha\!\in\!\{3,5,7,9\}$ (see sections 3.3 and 3.4) is tuned collectively for both V-DAB and T-DAB. We keep the kernel size $k$ of the interpolation functions $\phi(\cdot)$ and $\psi(\cdot)$ identical to $\alpha$ throughout the experiments. In all experiments, DeformTime has $2$ encoder layers (see section 3.5). For the ETT and weather tasks, we set $r\!=\!1$ in the first encoder layer and select from $r\!\in\!\{6,12,24\}$ time steps in the second encoder layer (see section 3.4). In addition, we select the segmentation length from $\ell\!\in\!\{6,12,24\}$ time steps (see section 3.3) collectively for both encoder layers. Given the more direct weekly temporal resolution in the reported ILI rates and the fact that web search activity trends do have weekly patterns, in the ILI forecasting task, we set $\ell\!=\!7$ time steps (days), and also set $r\!=\!1$ and $r\!=\!7$ time steps in the first and second encoder layer, respectively. The segmentation stride $s$ is kept the same as $\ell$ throughout the experiments.
Hyperparameters specific to forecasting tasks and optimisation settings.
The look-back window $L$ is set to $336$ time steps for the ETTh1, ETTh2, and weather forecasting tasks, based on Nie et al. (2023). For influenza forecasting, we set $L$ to $\{28,28,56,56\}$ days for forecasting horizons $H\!=\!\{7,14,21,28\}$ days ahead, respectively. Only for influenza forecasting, we also set the batch size and $d$ to $64$. For the other tasks, these become learnable parameters. The number of training epochs is set to a maximum of $100$ for ETTh1, ETTh2, and weather tasks and $50$ for the influenza forecasting task. Neural networks are optimised with Adam using mean squared error (MSE) loss on all outputs. As an exception, we use MAE loss in all models for the weather forecasting task as this gives better performance on the tuned baselines. Note that the output can be a time series (sequence) of one variable (the target) or more variables (target and exogenous) depending on the forecasting method.
[…]